# A quantum machine learning classifier to search for new physics

### v4

@@ Abstract
Due to the success of the Standard Model (SM), it is reasonable to anticipate that the signal of new physics (NP) beyond the SM is small. Consequently, future searches for NP and precision tests of the SM will require high luminosity collider experiments. Moreover, as precision tests advance, rare processes with many final-state particles require consideration which demand the analysis of a vast number of observables. The high luminosity produces a large amount of experimental data spanning a large observable space, posing a significant data-processing challenge. In recent years, quantum machine learning has emerged as a promising approach for processing large amounts of complex data on a quantum computer. In this study, we propose quantum searching neighbor (QSN) and variational QSN (VQSN) algorithms to search for NP. The QSN is a classification algorithm. The VQSN introduces variation to the QSN to process classical data. As applications, we apply the (V)QSN in the phenomenological study of the NP at the Large Hadron Collider and muon colliders. Examples are implemented on a real quantum hardware, which confirms reliable performance under noisy conditions. The results indicate that the VQSN demonstrates superior efficiency in the sense of computational complexity to a classical counterpart k-nearest neighbor algorithm, even when dealing with classical data.
@@ 1 Introduction
The standard model (SM) is supported by a substantial body of experimental evidence and can be concluded to describe and explain the majority of phenomena in particle physics, with a few rare exceptions. These exceptions include experimental results such as the neutrino mass Farzan:2017xzy ; Proceedings:2019qno ; Arguelles:2022tki , the WW boson mass CDF:2022hxs ; deBlas:2022hdk , the muon anomalous magnetic moment (g2g-2Muong-2:2021ojo ; Muong-2:2023cdq ; Aoyama:2020ynm , and other issues remain unresolved Crivellin:2023zui . Furthermore, the SM is unable to account for the existence of dark matter and the force of gravity. Consequently, there is a widely held belief in the existence of new physics (NP) beyond the SM, and the search for NP and precision tests of the SM have become a prominent area of interest within the high-energy physics (HEP) community Ellis:2012zz . Given the success of the SM, it is reasonable to posit that any NP signals that have yet to be discovered will be exceedingly weak. Therefore, the search for NP and the precision tests of the SM motivate higher luminosity collider experiments. The processing of large amounts of data is inherent to such experiments, and thus places demands on our ability to process large amounts of data efficiently.
@@
The application of machine learning (ML) algorithms represents a promising avenue for the efficient processing of data. Its applications have already been demonstrated in HEP Innocente:1992gq ; Albertsson:2018maf ; Guest:2018yhq ; Radovic:2018dip ; Baldi:2014kfa ; Ren:2017ymm ; Abdughani:2018wrw ; Ren:2019xhp ; Letizia:2022xbe ; DAgnolo:2019vbw ; DAgnolo:2018cun ; DeSimone:2018efk ; Lee:2018xtt . Of particular interest in the study of NP phenomenology is the employment of the anomaly detection (AD) algorithms, which have emerged as a prominent area of research in recent years MdAli:2020yzb ; Fol:2020tva ; Kasieczka:2021xcg ; CrispimRomao:2020ucc ; vanBeekveld:2020txa ; Kuusela:2011aa . Since ML methods often eliminate the need for detailed kinematic analysis, their benefits become increasingly evident as kinematics grow more complex. Consequently, the advantages will become increasingly apparent as further advances in the precision tests of the SM will inevitably involve the study of rare processes with larger numbers of particles in the final states.
@@
In the meantime, quantum computing represents another effective approach for processing large volumes of data. Despite quantum computing currently being in the noisy intermediate-scale quantum (NISQ) Chen:2021num ; Arute:2019zxq ; Preskill:2018jim era, there has been a notable increase in research activity within the HEP community related to this field Feynman:1981tf ; Roggero:2018hrn ; Roggero:2019myu ; Bauer:2022hpo ; Carena:2022kpg ; Gustafson:2022xdt ; Lamm:2024jnl ; Carena:2024dzu ; Atas:2021ext ; Li:2023vwx ; Cui:2019sfz ; Zou:2021pvl ; Georgescu:2013oza ; Lamm:2019uyc ; Li:2021kcs ; Echevarria:2020wct ; Perez-Salinas:2020nem ; Jordan:2011ci ; Mueller:2019qqj ; Zhu:2024own . Many ML algorithms can be implemented by quantum computing Biamonte:2016ugo ; qml2 ; Garcia:2022cqq . In conjunction with the advancement of quantum ML algorithms, there has been a growing interest in the deployment of quantum ML in the HEP phenomenological studies, encompassing techniques such as variational quantum classifier (VQC) Guan:2020bdl ; Wu:2020cye ; Grant:2018oml ; Terashi:2020wfi , quantum support vector machine (QSVM) Guan:2020bdl ; Wu:2021xsj ; Zhang:2023ykh ; Fadol:2022umw , and quantum kernel k-means (QKKM) Zhang:2024ebl . In this study, we propose a variational quantum searching neighbor (VQSN) algorithm for identifying NP signals.
@@
To illustrate our algorithm, the SM effective field theory (SMEFT) Weinberg:1979sa ; Grzadkowski:2010es ; Willenbrock:2014bja ; Masso:2014xra are used as the test cases. The SMEFT has recently been employed extensively in the search for NP. When the collision energy is insufficient to excite NP particles, the NP particles can be integrated out, and the NP effects become new interactions of known particles. These interactions are formally represented as higher-dimensional operators with Wilson coefficients suppressed by powers of a NP scale, denoted as $\Lambda$. The signals of the operators that are most likely to be found correspond to those with Wilson coefficients least suppressed by $\Lambda$. Instead of dealing with various NP models, the number of operators to be considered at a specific order of $\Lambda$ is finite. Moreover, if an operator is not found, all NP models contributing to this operator can be constrained. Therefore, the use of the SMEFT is in line with the theme of high efficiency in this work. The present study will investigate the utility of VQSN for the search for high operators at the large hadron collider (LHC) and future muon colliders. VQSN demonstrates superior performance compared to a classical k-nearest neighbor (KNN) algorithm, even when applied to classical datasets.
@@ 2.1 The quantum feature space
(V)QSN is a classification method. In the next section, we will explain why a classification method is used to search for NP, but in this section, we will focus on the classification method itself. The problem (V)QSN addresses is, if we know some points and their classifications, how should unknown points be classified? Fig. 1 provides an example. In Fig. 1, $32$ known points are classified into four categories (indicated by different colors). These $32$ points with known classification assignment constitute the training set and are generated using scikit-learn Pedregosa:2011ork , and rescaled to range $1<x_{i}<1$ and $1<y_{i}<1$ where $x_{i}$ and $y_{i}$ are the coordinates of the points. How should the unknown point (marked with an $\times$’) be classified? V(QSN) classifies points based on the distance between known points and unknown points. If an unknown point is closer to the known points of a certain class, it will be classified into that category.
@@ 2.2 The circuit to search for neighbors
As mentioned above, the problem becomes: given one $\phi$ and multiple $\psi_{j}$, how does one obtain the state $\langle\phi|\psi_{j}\rangle|j\rangle$. Assume $U_{|\phi\rangle}$ is the circuit that creates the quantum state $\phi$, i.e., $U_{|\phi\rangle}|0\rangle=|\phi\rangle$. For ease of description, we divide the registers into $a$ and $b$, where $a$ corresponds to the qubits storing $\psi_{j}$ and $b$ corresponds to the indices. For the state $|\psi_{j}\rangle_{a}|j\rangle_{b}$, if we apply the inverse of $U_{|\phi\rangle}$ to register $a$ and measuring the $a$ register yields the $|0\ldots 0\rangle_{a}$ state, then the state becomes $|0\ldots 0\rangle_{a}\langle\phi|\psi_{j}\rangle|j\rangle_{b}$. Note that the state corresponding to register $b$ at this point is precisely what we need. Denoting $U_{|\psi_{j}\rangle_{a}|j\rangle_{b}}$ as the circuit to create the state $|\psi_{j}\rangle_{a}|j\rangle_{b}$, i.e., $U_{|\psi_{j}\rangle_{a}|j\rangle_{b}}|0\rangle_{a}|0\rangle_{b}=|\psi_{j}\rangle_{a}|j\rangle_{b}$, the circuit is shown in Fig. 2. In this work, we use the uniform rotation gates Mottonen:2004vly to create the state $|\psi_{j}\rangle_{a}|j\rangle_{b}$
The success rate (denoted as $r$) to create the state $\langle\phi|\psi_{j}\rangle|j\rangle$ (the probabilities that $|0\ldots 0\rangle_{a}$ is measured) is $r=\sum\left|\langle\phi|\psi_{n}\rangle\right|^{2}/N$, where $N$ is the number of $\psi_{n}$ vectors. Without assumptions on the distribution of the vectors, there is no reason to prefer any result of the measurement on $a$ register. Therefore, if the vectors are $d$-dimensional complex vectors, $r\approx 1/d$, which is verified by using complex uniform random numbers. The success rate can be improved to $r\approx 1/\sqrt{d}$ by amplitude amplification, or the ‘QSearch’ algorithm  Brassard:2000xvp .
@@ 2.3 Fitting the training set
In this work, we use the ‘hardware efficient’ ansatz Kandala:2017vok ; Bravo-Prieto:2019kld with entanglement layers,
@@
where $\sigma^{q}_{y,z}$ are Pauli matrices acting on $q$ qubit, the product over $q$ is the product over all qubits, $c_{z}^{q_{1},q_{2}}$ is the controlled-Z gate acting on $q_{1,2}$ qubits with $c_{z}=diag\left(1,1,1,-1\right)$, and $\langle q_{1},q_{2}\rangle$ is the set of pairs of qubits depending on the type of entanglement layer, $l$ is the number of layers of the ansatz, $\{\alpha_{i}\}=\{\beta^{(q)}_{0},\rho^{(q)}_{0},\beta^{(q)}_{1},\rho^{(q)}_{1},\ldots\beta^{(q)}_{l},\rho^{(q)}_{l}\}$ is a set of trainable parameters. The number of parameters is $2(l+1)n_{q}$ where $n_{q}$ is the number of qubits. Two different entanglement layers are used in this work, which are the ‘shifted-circular-alternating’ (SCA) layer Sim:2019yyv and the ‘pairwise’ layer. An example of $U_{A}$ with the ‘SCA’ layer is shown in Fig. 3. The ‘pairwise’ layer is shown in Sec. 2.5.
@@ 2.4 Variation quantum searching neighbor algorithm
Since our goal is to classify, in the following, we assume that each $\psi$ vector carries a classification label and that the states are stored accordingly. Thus, the state to be create is $|\psi_{k,n}\rangle_{a}|k\rangle_{b_{1}}|n\rangle_{b_{2}}$, where $k$ indicates the class, the $b$ register is split to $b_{1}$ and $b_{2}$ for clarity. With this ordering, it becomes unnecessary to measure the entire $b$ register to find neighbors of $\phi$. Instead, only measuring $b_{1}$ reveals which class $\phi$’s neighbors belong to.
@@
    Using an ansatz circuit $|\Psi(\alpha_{i})\rangle$ to fit $|\psi_{k,n}\rangle_{a}|k\rangle_{b_{1}}|n\rangle_{b_{2}}$.
@@
In essence, the VQSN is a variant of the VQC algorithms. As a variant of VQC, given ansatze of comparable complexity their computational complexities are of the same order of magnitude. However, VQSN possesses two advantages. Firstly, conventional VQC can be seen as the quantum analogue of an artificial neural network, which is a ‘black box’ difficult to interpret. The reason is that VQC trains a parameterized quantum circuit to achieve the mapping $f(\psi_{sig})=1$ and $f(\psi_{bg})=0$, where $\phi_{sig}$ and $\phi_{bg}$ refer to the signal and background, respectively. This parameterized quantum circuit is constructed as a layered, interconnected ansatz, structurally resembling a neural network. The VQSN also trains a parameterized quantum circuit to realize the same discriminative effect. In contrast, the state to build in VQSN has a discernible geometric interpretation, and the target of the training is guaranteed to exist. Secondly, QSN is accompanied by a classical algorithm, KNN, which allows for a comparison between quantum and classical algorithms. This is clearer for a binary classification problem, in which QSN can be viewed as using $\sum w_{n}t_{n}/N$ to determine the classification of $\phi$ where $w_{n}=\left|\langle\phi|\psi_{n}\rangle\right|^{2}$, and $t_{n}=0,1$ is the class of $\psi_{n}$. Therefore, it uses a weighted average of the entire training dataset. KNN can also be viewed as a weighted average except that $\{w_{n}=1|\psi_{n}\in S\}$ and $\{w_{n}=0|\psi_{n}\notin S\}$, where $S$ is the set contains the $k$ nearest $\psi_{n}$ vectors to $\phi$.
Although our approach resembles KNN, it is fundamentally distinct. Unlike Refs. Lloyd:2013lby ; Wiebe:2014imz , we define distance through quantum overlap (same as Ref. basheer2021quantumknearestneighborsalgorithm ). This enables efficient identification of classification assignments via measurement. While QSN could extend to quantum KNN using the maxima finding algorithm from Ref. basheer2021quantumknearestneighborsalgorithm , VQSN deliberately eliminates this step to reduce quantum gate count and circuit depth. Furthermore, replacing data queries with an ansatz circuit significantly decreases gate requirements and depth, enabling execution of VQSN on a current NISQ hardware.

### v5

@@ Abstract
Due to the success of the Standard Model (SM), it is reasonable to anticipate that the signal of new physics (NP) beyond the SM is small. Consequently, future searches for NP and precision tests of the SM will require high luminosity collider experiments. Moreover, as precision tests advance, rare processes with many final-state particles require consideration which demands the analysis of a vast number of observables. The high luminosity produces a large amount of experimental data spanning a large observable space, posing a significant data-processing challenge. In recent years, quantum machine learning has emerged as a promising approach for processing large amounts of complex data on a quantum computer. In this study, we propose quantum searching neighbor (QSN) and variational QSN (VQSN) algorithms to search for NP. The QSN is a classification algorithm. The VQSN introduces variation to the QSN to process classical data. As applications, we apply the (V)QSN in the phenomenological study of the NP at the Large Hadron Collider and muon colliders. Examples are implemented on a real quantum hardware, which confirms reliable performance under noisy conditions. The results indicate that the VQSN demonstrates superior efficiency in the sense of computational complexity to a classical counterpart k-nearest neighbor algorithm, even when dealing with classical data.
@@ 1 Introduction
The standard model (SM) is supported by a substantial body of experimental evidence and can be concluded to describe and explain the majority of phenomena in particle physics, with a few rare exceptions. These exceptions include experimental results such as the neutrino mass Farzan and Tortola (2018); Bhupal Dev and others (2019); Argüelles and others (2023), the WW boson mass Aaltonen and others (2022); de Blas et al. (2022), the muon anomalous magnetic moment (g2g-2Abi and others (2021); Aguillard and others (2023); Aoyama and others (2020), and other issues remain unresolved Crivellin and Mellado (2023). Furthermore, the SM is unable to account for the existence of dark matter and the force of gravity. Consequently, there is a widely held belief in the existence of new physics (NP) beyond the SM, and the search for NP and precision tests of the SM have become a prominent area of interest within the high-energy physics (HEP) community Ellis (2012). Given the success of the SM, it is reasonable to posit that any NP signals that have yet to be discovered will be exceedingly weak. Therefore, the search for NP and the precision tests of the SM motivate higher luminosity collider experiments. The processing of large amounts of data is inherent to such experiments, and thus places demands on our ability to process large amounts of data efficiently.
@@
The application of machine learning (ML) algorithms represents a promising avenue for the efficient processing of data. Its applications have already been demonstrated in HEP Innocente et al. (1992); Albertsson and others (2018); Guest et al. (2018); Radovic et al. (2018); Baldi et al. (2014); Ren et al. (2019); Abdughani et al. (2019); Ren et al. (2020); Letizia et al. (2022); D’Agnolo et al. (2021); D’Agnolo and Wulzer (2019); De Simone and Jacques (2019); Lee et al. (2019). Of particular interest in the study of NP phenomenology is the employment of the anomaly detection (AD) algorithms, which have emerged as a prominent area of research in recent years Md Ali et al. (2020); Fol et al. (2020); Kasieczka and others (2021); Crispim Romão et al. (2021); van Beekveld et al. (2021); Kuusela et al. (2012). Since ML methods often eliminate the need for detailed kinematic analysis, their benefits become increasingly evident as kinematics grow more complex. Consequently, the advantages will become increasingly apparent as further advances in the precision tests of the SM will inevitably involve the study of rare processes with larger numbers of particles in the final states.
@@
In the meantime, quantum computing represents another effective approach for processing large volumes of data. Despite quantum computing currently being in the noisy intermediate-scale quantum (NISQ) Chen et al. (2021); Arute and others (2019); Preskill (2018) era, there has been a notable increase in research activity within the HEP community related to this field Feynman (1982); Roggero and Carlson (2019); Roggero et al. (2020); Bauer and others (2023); Carena et al. (2022); Gustafson et al. (2022); Lamm et al. (2024); Carena et al. (2024); Atas et al. (2021); Li et al. (2024); Cui et al. (2020); Zou et al. (2021); Georgescu et al. (2014); Lamm et al. (2020); Li et al. (2022); Echevarria et al. (2021); Pérez-Salinas et al. (2021); Jordan et al. (2014); Mueller et al. (2020); Zhu et al. (2024). Many ML algorithms can be implemented by quantum computing Biamonte et al. (2017); Schuld et al. (2015); García et al. (2022). In conjunction with the advancement of quantum ML algorithms, there has been a growing interest in the deployment of quantum ML in the HEP phenomenological studies, encompassing techniques such as variational quantum classifier (VQC) Guan et al. (2021); Wu and others (2021b); Grant et al. (2018); Terashi et al. (2021), quantum support vector machine (QSVM) Guan et al. (2021); Wu and others (2021a); Zhang et al. (2024b); Fadol et al. (2024), and quantum kernel k-means (QKKM) Zhang et al. (2024a). In this study, we propose a variational quantum searching neighbor (VQSN) algorithm for identifying NP signals.
@@
To illustrate our algorithm, the SM effective field theory (SMEFT) Weinberg (1979); Grzadkowski et al. (2010); Willenbrock and Zhang (2014); Masso (2014) are used as the test cases. The SMEFT has recently been employed extensively in the search for NP. When the collision energy is insufficient to excite NP particles, the NP particles can be integrated out, and the NP effects become new interactions of known particles. These interactions are formally represented as higher-dimensional operators with Wilson coefficients suppressed by powers of a NP scale, denoted as $\Lambda$. The signals of the operators that are most likely to be found correspond to those with Wilson coefficients least suppressed by $\Lambda$. Instead of dealing with various NP models, the number of operators to be considered at a specific order of $\Lambda$ is finite. Moreover, if an operator is not found, all NP models contributing to this operator can be constrained. Therefore, the use of the SMEFT is in line with the theme of high efficiency in this work. The present study will investigate the utility of VQSN for the search for high operators at the large hadron collider (LHC) and future muon colliders. VQSN demonstrates superior performance compared to a classical k-nearest neighbor (KNN) algorithm, even when applied to classical datasets.
@@ 2.1 The quantum feature space
(V)QSN is a classification method. In the next section, we will explain why a classification method is used to search for NP, but in this section, we will focus on the classification method itself. The problem (V)QSN addresses is, if we know some points and their classifications, how should unknown points be classified? Fig. 1 provides an example. In Fig. 1, $32$ known points are classified into four categories (indicated by different colors). These $32$ points with known classification assignment constitute the training set and are generated using scikit-learn Pedregosa and others (2011), and rescaled to range $1<x_{i}<1$ and $1<y_{i}<1$ where $x_{i}$ and $y_{i}$ are the coordinates of the points. How should the unknown point (marked with an $\times$’) be classified? V(QSN) classifies points based on the distance between known points and unknown points. If an unknown point is closer to the known points of a certain class, it will be classified into that category.
@@ 2.2 The circuit to search for neighbors
As mentioned above, the problem becomes: given one $\phi$ and multiple $\psi_{j}$, how does one obtain the state $\langle\phi|\psi_{j}\rangle|j\rangle$. Assume $U_{|\phi\rangle}$ is the circuit that creates the quantum state $\phi$, i.e., $U_{|\phi\rangle}|0\rangle=|\phi\rangle$. For ease of description, we divide the registers into $a$ and $b$, where $a$ corresponds to the qubits storing $\psi_{j}$ and $b$ corresponds to the indices. For the state $|\psi_{j}\rangle_{a}|j\rangle_{b}$, if we apply the inverse of $U_{|\phi\rangle}$ to register $a$ and measure the $a$ register yields the $|0\ldots 0\rangle_{a}$ state, then the state becomes $|0\ldots 0\rangle_{a}\langle\phi|\psi_{j}\rangle|j\rangle_{b}$. Note that the state corresponding to register $b$ at this point is precisely what we need. Denoting $U_{|\psi_{j}\rangle_{a}|j\rangle_{b}}$ as the circuit to create the state $|\psi_{j}\rangle_{a}|j\rangle_{b}$, i.e., $U_{|\psi_{j}\rangle_{a}|j\rangle_{b}}|0\rangle_{a}|0\rangle_{b}=|\psi_{j}\rangle_{a}|j\rangle_{b}$, the circuit is shown in Fig. 2. In this work, we use the uniform rotation gates Möttönen et al. (2004) to create the state $|\psi_{j}\rangle_{a}|j\rangle_{b}$
The success rate (denoted as $r$) to create the state $\langle\phi|\psi_{j}\rangle|j\rangle$ (the probabilities that $|0\ldots 0\rangle_{a}$ is measured) is $r=\sum\left|\langle\phi|\psi_{n}\rangle\right|^{2}/N$, where $N$ is the number of $\psi_{n}$ vectors. Without assumptions on the distribution of the vectors, there is no reason to prefer any result of the measurement on $a$ register. Therefore, if the vectors are $d$-dimensional complex vectors, $r\approx 1/d$, which is verified by using complex uniform random numbers. The success rate can be improved to $r\approx 1/\sqrt{d}$ by amplitude amplification, or the ‘QSearch’ algorithm  Brassard et al. (2000).
@@ 2.3 Fitting the training set
In this work, we use the ‘hardware efficient’ ansatz Kandala et al. (2017); Bravo-Prieto et al. (2023) with entanglement layers,
@@
where $\sigma^{q}_{y,z}$ are Pauli matrices acting on $q$ qubit, the product over $q$ is the product over all qubits, $c_{z}^{q_{1},q_{2}}$ is the controlled-Z gate acting on $q_{1,2}$ qubits with $c_{z}=diag\left(1,1,1,-1\right)$, and $\langle q_{1},q_{2}\rangle$ is the set of pairs of qubits depending on the type of entanglement layer, $l$ is the number of layers of the ansatz, $\{\alpha_{i}\}=\{\beta^{(q)}_{0},\rho^{(q)}_{0},\beta^{(q)}_{1},\rho^{(q)}_{1},\ldots\beta^{(q)}_{l},\rho^{(q)}_{l}\}$ is a set of trainable parameters. The number of parameters is $2(l+1)n_{q}$ where $n_{q}$ is the number of qubits. Two different entanglement layers are used in this work, which are the ‘shifted-circular-alternating’ (SCA) layer Sim et al. (2019) and the ‘pairwise’ layer. An example of $U_{A}$ with the ‘SCA’ layer is shown in Fig. 3. The ‘pairwise’ layer is shown in Sec. 2.5.
@@ 2.4 Variation quantum searching neighbor algorithm
Since our goal is to classify, in the following, we assume that each $\psi$ vector carries a classification label and that the states are stored accordingly. Thus, the state to be created is $|\psi_{k,n}\rangle_{a}|k\rangle_{b_{1}}|n\rangle_{b_{2}}$, where $k$ indicates the class, the $b$ register is split to $b_{1}$ and $b_{2}$ for clarity. With this ordering, it becomes unnecessary to measure the entire $b$ register to find neighbors of $\phi$. Instead, only measuring $b_{1}$ reveals which class $\phi$’s neighbors belong to.
@@
    Use an ansatz circuit $|\Psi(\alpha_{i})\rangle$ to fit $|\psi_{k,n}\rangle_{a}|k\rangle_{b_{1}}|n\rangle_{b_{2}}$.
@@
In essence, the VQSN is a variant of the VQC algorithms. As a variant of VQC, given ansatze of comparable complexity their computational complexities are of the same order of magnitude. However, VQSN possesses two advantages. Firstly, conventional VQC can be seen as the quantum analogue of an artificial neural network, which is a ‘black box’ difficult to interpret. The reason is that VQC trains a parameterized quantum circuit to achieve the mapping $f(\phi_{sig})=1$ and $f(\phi_{bg})=0$, where $\phi_{sig}$ and $\phi_{bg}$ refer to the signal and background, respectively. This parameterized quantum circuit is constructed as a layered, interconnected ansatz, structurally resembling a neural network. The VQSN also trains a parameterized quantum circuit to realize the same discriminative effect. In contrast, the state to build in VQSN has a discernible geometric interpretation, and the target of the training is guaranteed to exist. Secondly, QSN is accompanied by a classical algorithm, KNN, which allows for a comparison between quantum and classical algorithms. This is clearer for a binary classification problem, in which QSN can be viewed as using $\sum w_{n}t_{n}/N$ to determine the classification of $\phi$ where $w_{n}=\left|\langle\phi|\psi_{n}\rangle\right|^{2}$, and $t_{n}=0,1$ is the class of $\psi_{n}$. Therefore, it uses a weighted average of the entire training dataset. KNN can also be viewed as a weighted average except that $\{w_{n}=1|\psi_{n}\in S\}$ and $\{w_{n}=0|\psi_{n}\notin S\}$, where $S$ is the set contains the $k$ nearest $\psi_{n}$ vectors to $\phi$.
Although our approach resembles KNN, it is fundamentally distinct. Unlike Refs. Lloyd et al. (2013); Wiebe et al. (2015), we define distance through quantum overlap (same as Ref. Basheer et al. (2021)). This enables efficient identification of classification assignments via measurement. While QSN could extend to quantum KNN using the maxima finding algorithm from Ref. Basheer et al. (2021), VQSN deliberately eliminates this step to reduce quantum gate count and circuit depth. Furthermore, replacing data queries with an ansatz circuit significantly decreases gate requirements and depth, enabling execution of VQSN on a current NISQ hardware.